OpenStudy (anonymous):

How do you solve: 10/11x + 1/3 = 2/5 - 1/11x + 2/5

OpenStudy (anonymous):

The right side is just 4/5 - 1/11 x (by adding 2/5 and 2/5)

OpenStudy (anonymous):

Now we want to bring all the x's to one side of the equation, so add 1/11 x to both sides, you get 11/11 x + 1/3 = 4/5

OpenStudy (anonymous):

11/11 x is just 1x or x (becuase 11/11 = 1)

OpenStudy (anonymous):

So you have x + 1/3 = 4/5 Now subtract 1/3 from both sides, and what do you get, x = ?

OpenStudy (anonymous):

You can't subtract 1/3 from 4/5

OpenStudy (anonymous):

oh no?

OpenStudy (anonymous):

Yeah....

OpenStudy (anonymous):

Are you telling me that 4/5 - 1/3 cannot be performed?

OpenStudy (anonymous):

Well, you can perform a subtraction of 4/5 - 1/3. I will leave that to you.

OpenStudy (anonymous):

How

OpenStudy (anonymous):

Hint: Find the least common denominator, called LCD..which is 15 in this problem. Now you finish.

OpenStudy (anonymous):

x = 7/15 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh ok thanks

OpenStudy (anonymous):

welcome